# Cropping outperforms dropout as an augmentation strategy for training self-supervised text embeddings

### v1

@@ 3.2.3 Cropping-based fine-tuning is very fast
We further analyzed the performance improvement within the single fine-tuning epoch. We found that the MTEB block average was improved by around 15 percentage points within the first 100 fine-tuning batches (6 4006\,400 positive pairs) (Figure 1). Afterwards, the MTEB score plateaued for all modalities and did not improve any further, and fine-tuning for more than 1 epoch in the same dataset did not bring further improvements.
These 100 batches of fine-tuning took only $\sim$1 min of training time on a single GPU (NVIDIA RTX A6000). In comparison, the top-performing sentence transformer models are typically trained on large datasets, with substantial computational costs and training times. For example, the all-mpnet-base-v2 SBERT model was trained in a supervised way using over one billion text pairs. Even though its performance is higher (55.2%), we could bring the same base model (mpnet-base) close to SBERT’s performance on some of the tasks in a few minutes of self-supervised training using five orders of magnitude less data.
@@ 3.3 Representation of a dataset for analysis and visualization
Table 2: Representation of a given dataset. Score is kkNN accuracy of the mean pooling representation in percent ($k=10$ with Euclidean distance; we obtained similar values using cosine distance, see Table S2). We used a 9:1 train/test split for the kkNN classifier. Columns 1–2, 5–7: off-the-shelf models. Columns 3–4: MPNet fine-tuned on each dataset using cropping and dropout augmentations. Reported values should be interpreted with an error of up to $\pm$1%, corresponding to the binomial standard deviation $100\sqrt{p(1-p)/n}$ for test set size $n\approx 2000$ (smallest dataset) and accuracy $p=0.5$.
@@
3.3.1 Setup
We assessed the dataset representation quality through $k$-nearest-neighbor (kkNN) classification accuracy in the high-dimensional ($d=768$) embedding space. This metric is computed by classifying each point according to the majority class among its $k$ nearest neighbors, and then comparing predicted class against the true class. It is a measure of local coherence: it is high if each point’s nearest neighbors belong to the same class as the point itself. This metric is particularly relevant for data exploration applications, e.g. visualization or clustering, as many unsupervised learning algorithms for these tasks rely on the kkNN graph of the data. While this metric depends on the quality of the kkNN graph similarly to retrieval or reranking tasks from MTEB, it is simpler to evaluate as it requires only class labels rather than annotated samples or ranked neighbors.
Figure 2: $t$-SNE visualizations of MPNet, SBERT, and cropping-based fine-tuned model embeddings of different datasets. Color corresponds to class labels. Numbers show kkNN accuracy in 768D embedding space. We used openTSNE with default parameters (Poličar et al., 2024).
We performed our fine-tuning experiments on six datasets: the arXiv, bioRxiv, medRxiv, Reddit, and StackExchange datasets from the P2P clustering tasks of the MTEB (Muennighoff et al., 2023), and the ICLR dataset (González-Márquez & Kobak, 2024). The datasets differed in the number of samples (18–733 thousand) and classes (26–610; Table S4). Four of them comprised scientific abstracts from different disciplines, and the other two consisted of internet posts.
@@ 3.3.2 Results
We found that, on average across datasets, the base MPNet produced representations with low accuracy (46.5%, Table 2) and almost no semantic structure visible in 2D visualizations using $t$-SNE (van der Maaten & Hinton, 2008) (Figure 2). Fine-tuning MPNet for one epoch for each dataset increased the performance to 55.8% on average, which was still below SBERT with 59.5%. For some datasets this improvement was particularly large; e.g., the representation of the ICLR dataset improved by over 20 percentage points.
@@ 3.3.3 Sentence and domain adaptation during fine-tuning
When fine-tuning a token-level base model with a sentence-level contrastive loss on a specific dataset, two mechanisms can contribute to the performance improvement: the model adapting to represent sentences (“sentence adaptation”) and the model adapting to the domain of the training data (“domain adaptation”). To disentangle the contributions of these two potential mechanisms, we performed self-supervised fine-tuning on one dataset and evaluated the model’s performance on a dataset from a different domain. We used three MTEB clustering P2P datasets from the previous section (arXiv, bioRxiv, and Reddit) as training datasets, and ICLR as the evaluation dataset. For comparable training conditions, we used dataset subsets equal to ICLR’s size. We fine-tuned the base model separately on each dataset and measured the kkNN accuracy on the ICLR dataset. We also fine-tuned and evaluated directly on ICLR for comparison with the setting when both adaptation mechanisms are present.
Figure 3: Domain vs. sentence adaptation. kkNN accuracy on the ICLR dataset for MPNet fine-tuned separately on four different datasets (arXiv, bioRxiv, Reddit, ICLR). Most of the improvement happens within the first 100 batches.
@@
3.3.4 Self-supervised training without pre-training
To determine whether the token-level pre-training was necessary to achieve good sentence representations of a given dataset, we performed cropping-based contrastive training of the bert-base architecture from scratch, without using pre-trained MPNet weights (Table S5, column 10). Here, the performance did not saturate after one epoch, so we continued training for 10 epochs. On average across datasets, the resulting performance was 2.6 percentage points below the one we obtained from fine-tuning the pre-trained MPNet, and for some of the datasets there was no noticeable performance difference at all. However, when training the model on ICLR and evaluating on MTEB (setup from Section 3.2) the results were much worse compared to using the pre-trained MPNet (18.9 p.p. difference).
As an additional ablation, we performed the same cropping-based contrastive training of a bare, randomly initialized, embedding layer. This is a direct token embedding model without any transformer architecture whatsoever. Training it for 10 epochs, we obtained dataset embeddings that on average were 3.5 percentage points below the ones obtained from fine-tuning the pre-trained MPNet (Table S5, column 11). Again, when trained on ICLR and evaluated on MTEB, the results were very poor, with the block average much worse compared to using the pre-trained MPNet (13.4 p.p. difference).
These results highlight that even the simple token embedding layer can achieve a reasonable embedding of a given dataset, when trained with cropping augmentations (dropout augmentation worked much worse; Table S5). However, the transformer architecture and pre-training knowledge are clearly necessary when generalizing to other tasks and domains.
@@ 4 Representation quality across layers
Figure 4: Representation quality across layers. (a) kkNN accuracy after fine-tuning MPNet with different number of initial layers frozen. The embedding layer was frozen in all settings. Zero unfrozen layers corresponds to no fine-tuning. (b) kkNN accuracy after each layer for MPNet before and after fine-tuning, for SBERT, for randomly initialized model, and for model trained with cropping augmentations from scratch. Layer 0 corresponds to the embedding layer.
@@
To investigate whether fine-tuning the full MPNet model was needed to obtain high-quality sentence representations, we performed cropping-based fine-tuning on the ICLR dataset while freezing the embedding layer and various numbers of initial layers. We observed that the performance rapidly improved with the number of unfrozen layers, and fine-tuning only the last two out of 12 layers for one epoch was sufficient to reach almost the same value of kkNN classification accuracy as fine-tuning the full model (Figure 4a). Unfreezing additional layers led only to minor further improvements.
@@
When all layers were unfrozen, the last few layers underwent the largest change during fine-tuning, while the early layers barely changed, in agreement with previous findings in supervised setting (Merchant et al., 2020; Mosbach et al., 2020). To quantify this, we measured the representation quality after each hidden layer before and after fine-tuning (Figure 4b). The gap between them was close to zero for early layers and increased towards the last layers. We observed the same effect when fine-tuning MPNet on other datasets (Figure S2).
@@
Intriguingly, the representation quality across layers in our fine-tuned model as well as in out-of-the-box SBERT formed a U-shaped curve (Figure 4b): before fine-tuning the embedding layer representation had the highest accuracy, and after fine-tuning it was surpassed by only the last two layers. Across other datasets, the shape was different and not always U-shaped (Figure S2), but fine-tuned models always exhibited a steep rise in performance towards layer 12. The randomly initialized models did not exhibit this shape: after SSL training, the performance monotonically increased and plateaued half-way through the layers (Figure 4b).
Consistently across all datasets and fine-tuned models, the last layer always gave the best representation (Figures 4b and S2). This differs from what has been observed in computer vision, where the top performance after SSL training typically occurs in one of the hidden layers. Indeed, a common practice in computer vision is to have several fully-connected layers (projection head) between the output representation and the contrastive loss (Chen et al., 2020), which are discarded after SSL training (guillotine regularization) (Bordes et al., 2023). We experimented with adding a one-hidden-layer ($768\to 512\to 128$) projection head after the average pooling, but this did not consistently affect the representation quality, in agreement with Figure 4b.
@@ 5 Discussion
We showed that self-supervised fine-tuning is a powerful strategy for producing high-quality text embeddings with minimal training on in-domain data. To this end, we systematically compared different self-supervised augmentation techniques under the exact same training setup and showed that cropping augmentations were much better than dropout augmentations in all evaluations. This finding is noteworthy because dropout augmentations of SimCSE have been claimed to be superior (Gao et al., 2021) and are one of the most well-cited SSL approaches in the literature on text embeddings.
While we showed that the cropping augmentation can greatly enhance text representations of multi-sentence texts, we believe that there is still a large room for improvement. The lesson from computer vision (Balestriero et al., 2023) as well as from our work is that good data augmentations are crucial for the success of self-supervised learning. Combining cropping augmentation with more powerful semantic augmentations such as reformulations using generative language models (Jiang et al., 2022; Wang & Dou, 2023; Abaskohi et al., 2023) can offer an interesting avenue for future research.
Despite self-supervised fine-tuning substantially improving text embedding quality, there was still a gap until SOTA supervised models. It is unclear whether this gap stems from supervision itself or from the much larger and more diverse datasets on which SOTA supervised models are trained. Such models typically leverage diverse data spanning multiple domains, minimizing out-of-domain scenarios — in contrast to our limited self-supervised training on a single dataset. It remains an open question whether this gap can be bridged with extensive self-supervised training on diverse data or whether supervision is a key element to achieve better performance.
@@ Acknowledgments
The authors would like to thank Carsten Eickhoff for feedback and discussions. This work was funded by the Deutsche Forschungsgemeinschaft (KO6282/2-1) and by the Gemeinnützige Hertie-Stiftung. Philipp Berens and Dmitry Kobak are members of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Rita González Márquez.
@@ A.1 Augmentation and hyperparameter choices for self-supervised fine-tuning
To select the optimal hyperparameters for our self-supervised fine-tuning, we performed a detailed hyperparameter study using the ICLR dataset. We ran the fine-tuning of MPNet for one epoch, and assessed the final kkNN classification accuracy (Figure S1). In each experiment, all other parameters were kept at their default values described in Section 3.1.
@@
We compared four different pooling strategies for forming the final sentence representation: average pooling, the classification token [CLS], the separation token [SEP] (appended at the end of each input text), and the seventh token (as an example of an arbitrary token number). We obtained the best results using the average pooling and [SEP] token, with the other two options performing less well (Figure S1a).
When evaluating off-the-shelf models, we always used the mean pooling. On the ICLR dataset, some off-the-shelf models showed slightly higher kkNN accuracy in the [SEP] token representation than using the [CLS] token or average pooling, but the difference was small (Table S6). It has recently been shown in a computer vision setting that additional tokens can be used by the transformer model as ‘registers’ to store high-level features (Darcet et al., 2023). Our results suggest that the same can happen with language models, since the [SEP] token often serves as a good sentence representation despite not being explicitly used for training.
@@
Our data augmentation consisted of ‘cropping out’ $t$ consecutive sentences. We varied the number of consecutive sentences (decreasing the batch size accordingly, to make it fit into the GPU memory) and found that the performance generally decreased with $t$, with the optimal number being $t=2$ (Figure S1b). Note that in our sampling it was possible for the positive pair of text chunks to overlap (but not to coincide exactly).

### v2

@@ 3.2.3 Cropping-based fine-tuning is very fast
We further analyzed the performance improvement within the single fine-tuning epoch. We found that the MTEB block average was improved by around 15 percentage points within the first 100 fine-tuning batches (6 4006\,400 positive pairs) (Figure 1). Afterwards, the MTEB score plateaued for all modalities and did not improve any further, and fine-tuning for more than 1 epoch in the same dataset did not bring further improvements. As a caveat, it is possible that one would observe further slower gains in performance with larger amount of training data and longer training duration.
Figure 1: MTEB score during fine-tuning. Improvement on block mean MTEB score and individual task modalities within the first fine-tuning epoch on ICLR dataset.
These 100 batches of fine-tuning took only $\sim$1 min of training time on a single GPU (NVIDIA RTX A6000). In comparison, the top-performing sentence transformer models are typically trained on large datasets, with substantial computational costs and training times. For example, the all-mpnet-base-v2 SBERT model was trained in a supervised way using over one billion text pairs. Even though its performance is higher (58.8%), we could bring the same base model (mpnet-base) close to SBERT’s performance on some of the tasks in a few minutes of self-supervised training using five orders of magnitude less data.
@@ 3.3 Representation of a dataset for analysis and visualization
3.3.1 Setup
We assessed the dataset representation quality through $k$-nearest-neighbor (kkNN) classification accuracy in the high-dimensional ($d=768$) embedding space. This metric is computed by classifying each point according to the majority class among its $k$ nearest neighbors, and then comparing predicted class against the true class. We used $k=10$ with Euclidean distance and non-normalized embeddings, but we obtained similar results using cosine distance (which is equivalent to using normalized embeddings; Table S3). It is a measure of local coherence: it is high if each point’s nearest neighbors belong to the same class as the point itself. This metric is particularly relevant for data exploration applications, e.g. visualization or clustering, as many unsupervised learning algorithms for these tasks rely on the kkNN graph of the data. While this metric depends on the quality of the kkNN graph similarly to retrieval or reranking tasks from MTEB, it is simpler to evaluate as it requires only class labels rather than annotated samples or ranked neighbors.
We performed our fine-tuning experiments on six datasets: the arXiv, bioRxiv, medRxiv, Reddit, and StackExchange datasets from the P2P clustering tasks of the MTEB (Muennighoff et al., 2023), and the ICLR dataset (González-Márquez and Kobak, 2024). The datasets differed in the number of samples (18–733 thousand) and classes (26–610; Table S5). Four of them comprised scientific abstracts from different disciplines, and the other two consisted of internet posts.
Table 2: Representation quality of a given dataset. kkNN accuracy of the mean-pooling representation in percent ($k=10$). We used a 9:1 train/test split for the kkNN classifier. Columns 1–2, 5–7: off-the-shelf models. Columns 3–4: MPNet fine-tuned on each dataset using cropping and dropout augmentations. Reported values should be interpreted with an error of up to $\pm$1%, corresponding to the binomial standard deviation $100\sqrt{p(1-p)/n}$ for test set size $n\approx 2000$ (smallest dataset) and accuracy $p=0.5$.
@@ 3.3.2 Results
Figure 2: Dataset visualizations. $t$-SNE visualizations of MPNet, SBERT, and cropping-based fine-tuned MPNet embeddings of different datasets. Color corresponds to class labels. Numbers show kkNN accuracy in 768D embedding space. We used openTSNE with default parameters (Poličar et al., 2024). The fine-tuned model was always fine-tuned on the same dataset, as in Table 2.
We found that, on average across datasets, the base MPNet produced representations with low accuracy (46.5%, Table 2) and almost no semantic structure visible in 2D visualizations using $t$-SNE (van der Maaten and Hinton, 2008) (Figure 2). Fine-tuning MPNet for one epoch for each dataset increased the performance to 55.8% on average, which was still below SBERT with 59.5%. For some datasets this improvement was particularly large; e.g. the representation of the ICLR dataset improved by over 20 percentage points.
@@ 3.3.3 Sentence and domain adaptation during fine-tuning
When fine-tuning a base model pre-trained on the token level with a sentence-level contrastive loss on a specific dataset, two mechanisms can contribute to the performance improvement: the model adapting to represent sentences (“sentence adaptation”) and the model adapting to the domain of the training data (“domain adaptation”). To disentangle the contributions of these two potential mechanisms, we performed self-supervised fine-tuning on one dataset and evaluated the model’s performance on a dataset from a different domain. We used three MTEB clustering P2P datasets from the previous section (arXiv, bioRxiv, and Reddit) as training datasets, and ICLR as the evaluation dataset. For comparable training conditions, we used dataset subsets equal to ICLR’s size. We fine-tuned the base model separately on each dataset and measured the kkNN accuracy on the ICLR dataset. We also fine-tuned and evaluated directly on ICLR for comparison with the setting when both adaptation mechanisms are present.
Figure 3: Sentence vs. domain adaptation. kkNN accuracy on the ICLR dataset for MPNet fine-tuned separately on four different datasets (arXiv, bioRxiv, Reddit, ICLR).
@@
3.4 Self-supervised training without pre-training
To determine whether the token-level pre-training was necessary to achieve good sentence representations of a given dataset, we performed cropping-based contrastive training of the bert-base architecture from scratch, without using pre-trained MPNet weights (Table S6, column 10). Here, the performance did not saturate after one epoch, so we continued training for 10 epochs. On average across datasets, the resulting kkNN accuracy was only 2.6 percentage points below the one we obtained from fine-tuning the pre-trained MPNet, and for some of the datasets there was no noticeable performance difference at all. However, when training the model on ICLR and evaluating on MTEB (setup from Section 3.2 but without classification tasks) the results were much worse compared to using the pre-trained MPNet for fine-tuning (18.9 p.p. difference).
As an additional ablation, we performed the same cropping-based contrastive training of a bare, randomly initialized, embedding layer. This is a direct token embedding model without any transformer architecture whatsoever. Training it for 10 epochs, we obtained dataset embeddings that on average were 3.5 percentage points below the ones obtained from fine-tuning the pre-trained MPNet in kkNN accuracy (Table S6, column 11). Again, when trained on ICLR and evaluated on MTEB (without classification tasks; non-normalized embeddings), the results were very poor, with a much worse block average compared to the full pre-trained MPNet with additional cropping fine-tuning (13.4 p.p. difference).
These results highlight that even the simple token embedding layer can achieve reasonable embeddings of a given dataset when trained with cropping augmentations (dropout augmentation worked much worse; Table S6). However, the transformer architecture and pre-training knowledge are clearly necessary when generalizing to other tasks and domains.
@@ 4 Representation quality across layers
Figure 4: Representation quality across layers. (a) kkNN accuracy after fine-tuning MPNet with different number of initial layers frozen. The embedding layer was frozen in all settings. Zero unfrozen layers corresponds to no fine-tuning. (b) kkNN accuracy after each layer for MPNet before and after fine-tuning all layers, for SBERT, for a randomly initialized model with and without training with cropping augmentations. Layer 0 corresponds to the embedding layer. (c) MTEB block average score after each layer for MPNet before and after fine-tuning with both kinds of augmentations, and for SBERT. Layer 0 corresponds to the embedding layer. Here the text embeddings were not normalized for the evaluation, unlike in Section 3.2, therefore exact values are slightly different from Table 1.
@@
To investigate whether fine-tuning the entire MPNet model was necessary to obtain high-quality sentence representations, we performed cropping-based fine-tuning on the ICLR dataset while freezing the embedding layer and various numbers of initial layers. We observed that the performance rapidly improved with the number of unfrozen layers, and fine-tuning only the last 2 out of 12 layers for one epoch was sufficient to reach almost the same value of kkNN classification accuracy as fine-tuning the full model (Figure 4a). Unfreezing additional layers led only to minor further improvements.
@@
When all layers were unfrozen, the last few layers underwent the largest change during fine-tuning, while the early layers barely changed, in agreement with previous findings in the supervised setting (Merchant et al., 2020; Mosbach et al., 2020). To quantify this, we measured the representation quality after each hidden layer before and after fine-tuning the full model (Figure 4b–c). The gap between them was close to zero for early layers and increased towards the last layers (for both kinds of augmentations: dropout and crops). We observed the same effect when fine-tuning MPNet on other datasets and using the kkNN evaluation (Figure S2).
@@
Intriguingly, the representation quality across layers in our fine-tuned model as well as in out-of-the-box SBERT formed a U-shaped curve (Figure 4b–c): before fine-tuning the embedding layer representation had the highest kkNN accuracy and almost the highest MTEB score, and after fine-tuning it was surpassed by only the last layers. Across other datasets, the shape was different and not always U-shaped (Figure S2), but fine-tuned models always exhibited a steep rise in performance towards layer 12. The randomly initialized models did not exhibit this shape: after SSL training, the performance monotonically increased and plateaued half-way through the layers (Figure 4b).
It has been argued that misalignment between the training objective and the downstream task can lead to representation quality peaking in the middle layers (Bordes et al., 2023). This does not explain why the representation quality of non-fine-tuned MPNet sometimes was the highest in the embedding layer, but prior work found similar results when evaluating BERT and some other language models on MTEB (Skean et al., 2025). However, none of these models had been fine-tuned for text embeddings, so evaluating their text-embedding performance is not very informative. Our results extend those by Skean et al. (2025) and show that, once fine-tuned for text representation, models’ representation quality does increase towards the last layers.
A common practice in computer vision is to have several fully-connected layers (projection head) between the last layer and the contrastive loss (Chen et al., 2020), which are discarded after SSL training (guillotine regularization), to correct for the misalignment between training and evaluation (Bordes et al., 2023). Our results show that this was not necessary when using the cropping augmentation, as the representation quality peaked in the last layer. However, with dropout augmentation and MTEB evaluation, the representation quality peaked in the penultimate layer (Figure 4c). Importantly, it was still lower than with the cropping augmentation. When analyzing MTEB task categories separately, we saw a similar effect with cropping augmentation and some of the MTEB tasks (classification and STS), where representation quality peaked in the penultimate layer (Figure S3).
5 Limitations
Our main results focus on two augmentation techniques — cropping and dropout — due to their popularity in the text-embedding literature. We explored additional strategies in pilot experiments (random masking of a fraction of the words, and a combination of that with cropping; Section A.1), but these performed worse compared to cropping and we therefore excluded them from the full analysis. An interesting direction for future work would be to compare cropping augmentation to LLM-generated contrastive pairs based on rephrasing or reformulation of the anchor (Jiang et al., 2022; Wang and Dou, 2023; Abaskohi et al., 2023).
Another direction of future research is to investigate the effect of using hard negatives during training. Hard negatives are negative samples that are semantically or structurally similar to the anchor. They can be identified using external sources of information (such as citations; Cohan et al., 2020; Ostendorff et al., 2022), i.e. in a supervised manner. Recent works have also investigated self-supervised ways of mining hard negatives (Xiong et al., 2021; Liu et al., 2024; Li et al., 2024) and argued that this can lead to improvement performance.
Our comparison of the two augmentation techniques on MTEB tasks relies on models fine-tuned on the ICLR dataset (Section 3.2), representing an out-of-domain evaluation scenario. While we studied performance differences across multiple training datasets in the in-domain setting (kkNN accuracy, Section 3.3), we did not extend this to the MTEB evaluation. The main reason was that we wanted to avoid any train-test overlap and chose the ICLR dataset for training as it is not part of the MTEB evaluations.
@@ 6 Discussion
We showed that self-supervised fine-tuning is a powerful strategy for producing high-quality text embeddings with minimal training on in-domain data. To this end, we systematically compared different self-supervised augmentation techniques under the exact same training setup and showed that cropping augmentations were much better than dropout augmentations in all evaluation modalities. This finding is noteworthy because dropout augmentations are one of the most well-cited SSL approaches in the literature on text embeddings (Gao et al., 2021) and continue to be sometimes used for training modern text embedding models (BehnamGhader et al., 2024).
Despite self-supervised fine-tuning substantially improving text embedding quality, there was still a gap compared to the supervised SOTA models. It is unclear whether this gap stems from supervision itself or from the much larger and more diverse datasets on which SOTA supervised models are trained. Such models typically leverage diverse data spanning multiple domains, minimizing out-of-domain scenarios — in contrast to our limited self-supervised training on a single dataset. It remains an open question whether this gap can be bridged with extensive self-supervised training on diverse data or whether supervision is a key element to achieve better performance.
@@ Acknowledgments
The authors would like to thank Carsten Eickhoff for feedback and discussions. This work was funded by the Deutsche Forschungsgemeinschaft (KO6282/2-1) and by the Gemeinnützige Hertie-Stiftung. Philipp Berens and Dmitry Kobak are members of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 2064, ref 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Rita González Márquez.
@@ A.1 Augmentation and hyperparameter choices for self-supervised fine-tuning
To select the optimal hyperparameters for our self-supervised fine-tuning, we performed a detailed hyperparameter analysis using the ICLR dataset. We ran the fine-tuning of MPNet for one epoch, and assessed the final kkNN classification accuracy (Figure S1). In each experiment, all other parameters were kept at their default values described in Section 3.1.
@@
We compared four different pooling strategies for forming the final sentence representation: average pooling, the classification token [CLS], the separation token [SEP] (appended at the end of each input text; similar to [EOS] token), and the seventh token (as an example of an arbitrary token). We obtained the best results using the average pooling and [SEP] token, with the other two options performing less well (Figure S1a).
When evaluating off-the-shelf models, we always used the mean pooling. On the ICLR dataset, some off-the-shelf models showed slightly higher kkNN accuracy in the [SEP] token representation than using the [CLS] token or average pooling, but the difference was small (Table S7). It has recently been shown in a computer vision setting that additional tokens can be used by the transformer model as ‘registers’ to store high-level features (Darcet et al., 2024). Our results suggest that the same can happen with language models, since the [SEP] token often serves as a good sentence representation despite not being explicitly used for training.
@@
Our data augmentation consisted of ‘cropping out’ $t$ consecutive sentences. We used periods (‘.’) as sentence delimitators, and discarded sentences shorter than 100 characters and longer than 250 characters. We varied the number of consecutive sentences (decreasing the batch size accordingly, to make it fit into the GPU memory) and found that the performance generally decreased with $t$, with the optimal number being $t=2$ (Figure S1b). Note that in our sampling it was possible for the positive pair of text chunks to overlap (but not to coincide exactly).